Find the present value and the future value of the following four year mixed stream cash flow at 12% p.a.
Year 1 cash flow is 40000, year 2 cash flow is 800000, year 3 cash flow is 40000 and year 4 cash flow is 50000
1
Expert's answer
2020-12-14T19:52:17-0500

"FV=(40\cdot1.12^4+800\cdot1.12^3+40\cdot1.12^2+50\cdot1.12)\cdot1000=1293059."

"PV=\frac{1293059}{1.12^4}=821762."